OpenStudy (anonymous):

ok here is the problem. 3+sqrt(z-10)=sqrt(x+5). I know how to solve it, and the steps, but don't understand the steps if that makes sense. I know the next step is 9+3sqrt(z-10)+3sqrt(z-10)+z-10=z+5, but do not understand how you get to that point. Can someone please give me a step by step on how to get it?

OpenStudy (amistre64):

the only step that may seem odd is squaring everthing, but thats only because thats the undo process of the square root

OpenStudy (anonymous):

i know you have to square both sides, which to me would be 9 + z -10=z+5

OpenStudy (amistre64):

this mouse is awful

OpenStudy (amistre64):

3+sqrt(z-10)=sqrt(x+5) ; ^2 9 +6sqrt(z-10) +(z-10) = (x+5)

OpenStudy (amistre64):

or if we do the multiplication out like normally 3+sqrt(z-10) 3+sqrt(z-10) ------------ 9 + 3sqrt(z-10) 3sqrt(z-10) + (z-10) ---------------------- 9 +6sqrt(z-10) + (z-10)

OpenStudy (amistre64):

a+b a+b ------- a^2 + ab ab + b^2 -------------- a^2 +2ab +b^2

OpenStudy (anonymous):

great! I didn't do the squaring right for the first 1/2 of the problem, i forgot about FOIL when doing it lol. Thanks!